# MUSE: Flexible Voiceprint Receptive Fields and Multi-Path Fusion Enhanced Taylor Transformer for U-Net-based Speech Enhancement

### v1

@@ 1 Introduction
For previous methods, a higher number of channels (64) is often required to achieve peak performance. The popular Two-Stage (TS) [7] structural network in speech enhancement demonstrates strong competitiveness. The TS structure typically performs downsampling only once along the temporal dimension due to the large size of the feature maps. This implies that such structures require greater memory and computational resources, particularly as the length of the speech increases, necessitating significant GPU memory during both training and inference stages. Training or deploying such speech enhancement models on devices with limited GPU memory presents evident challenges. To mitigate GPU memory consumption for the same length of speech, the simplest approach is to reduce the number of channels to decrease model size. However, for speech enhancement tasks, the model needs to map speech features to a high-dimensional space; insufficient channel numbers often result in decreased ability to discriminate noise features from speech features, potentially leading to significant performance degradation. We retrained the state-of-the-art (SOTA) model MP-SEnet [11] with 16 channels as a baseline, subsequently referred to as MP-SEnet-16. Experimental results indicate that while MP-SEnet-16 significantly reduces GPU memory requirements and parameter count compared to its 64 channels versions, the model’s performance also significantly deteriorates. Therefore, for scenarios with low channel counts, we consider continuing to use the TS structure as suboptimal. Further more, for voiceprint information in speech, the feature shape often resembles an elongated crescent, contrasting with normal conv kernels that are fixed in a square shape. Hence, learning features from those specific shaped voiceprint becomes challenging and inefficient.
@@ 2.2 Deformable Embedding
In the realm of acoustic signal processing, sound frequencies and features exhibit significant variability. Voiceprints typically possess unique shapes, thus, convolutional receptive fields may capture specific voiceprint features inadequately. In pursuit of a more adaptable receptive field configuration, akin to strategies employed in the field of image processing, we introduce depthwise separable and deformable convolutions (DSDCN) [15]. Through the application of depthwise convolution and pointwise convolution, we mitigated computational complexity and reduced parameter count, consequently enhancing the efficiency of speech signal processing. We have incorporated Deformable Embedding at the head of each U-net encoder-decoder, as illustrated by the MET Transformer block in Figure 2. DSDCN concurrently possesses the capability to learn intricate details of voiceprint information at fine scales and large-scale voiceprint features. This implies that DSDCN not only has the capability to capture the shape of voiceprint features but also exhibits the ability to expand the receptive field, similar to the behavior of dilated convolutions. Hardswish activation is employed to harness heightened non-linear characteristics, enabling the network to extract more intricate acoustic features. It is noteworthy that the application of DSDCN is avoided in the input encoder and final magnitude-phase decoder of the network. This strategic exclusion arises from the inherent flexibility of deformable convolutions in achieving adaptable receptive fields through feature offsets, as the use of DCN in the proximate stages of input and output results in diminished performance.
@@ 2.4 Dense Convolution Codec
In the input-output encoder, we borrowed the Codec from MP-SEnet [11]. The expanded Dense-Net structure comprises four convolutional blocks with densely connected residual links [23], each block having dilation factors set to ${1,2,4,8}$. In dilated convolutions, our primary objective is to augment the receptive field while preserving the convolutional kernel and layer counts, ensuring the efficient capture of globally and intricately featured information.
@@ 3.2 Model setup
Throughout the training process, speech data was uniformly segmented into 30700 points, employing an FFT size of 510, a window length of 510, a hop length of 100, and a sampling rate of 16000  [^1] https://github.com/huaidanquede/MUSE-Speech-Enhancement. The training configuration encompassed a batch size of 2, a learning rate set to 0.0005 with a decay factor of 0.99, dense channels is 16, all the models were trained using the AdamW [30] opti mizer until 100 epochs. During training, a single 8GB RTX 3070ti GPU was utilized.

### v2

@@ 1 Introduction
For previous methods, a higher number of channels (64) is often required to achieve peak performance. The popular Two-Stage (TS) [7] structural network in speech enhancement demonstrates strong competitiveness. The TS structure typically performs downsampling only once along the frequency dimension due to the large size of the feature maps. This implies that such structures require greater memory and computational resources, particularly as the length of the speech increases, necessitating significant GPU memory during both training and inference stages. Training or deploying such speech enhancement models on devices with limited GPU memory presents evident challenges. To mitigate GPU memory consumption for the same length of speech, the simplest approach is to reduce the number of channels to decrease model size. However, for speech enhancement tasks, the model needs to map speech features to a high-dimensional space; insufficient channel numbers often result in decreased ability to discriminate noise features from speech features, potentially leading to significant performance degradation. We retrained the state-of-the-art (SOTA) model MP-SEnet [11] with 16 channels as a baseline, subsequently referred to as MP-SEnet-16. Experimental results indicate that while MP-SEnet-16 significantly reduces GPU memory requirements and parameter count compared to its 64 channels versions, the model’s performance also significantly deteriorates. Therefore, for scenarios with low channel counts, we consider continuing to use the TS structure as suboptimal. Further more, for voiceprint information in speech, the feature shape often resembles an elongated crescent, contrasting with normal conv kernels that are fixed in a square shape. Hence, learning features from those specific shaped voiceprint becomes challenging and inefficient.
@@ 2.2 Deformable Embedding
In the realm of acoustic signal processing, sound frequencies and features exhibit significant variability. Voiceprints typically possess unique shapes, thus, convolutional receptive fields may capture specific voiceprint features inadequately. In pursuit of a more adaptable receptive field configuration, akin to strategies employed in the field of image processing, we introduce depthwise separable and deformable convolutions (DSDCN) [15]. Through the application of depthwise convolution and pointwise convolution, we mitigated computational complexity and reduced parameter count, consequently enhancing the efficiency of speech signal processing. We have incorporated Deformable Embedding at the head of each U-Net encoder-decoder, as illustrated by the MET Transformer block in Figure 2. DSDCN concurrently possesses the capability to learn intricate details of voiceprint information at fine scales and large-scale voiceprint features. This implies that DSDCN not only has the capability to capture the shape of voiceprint features but also exhibits the ability to expand the receptive field, similar to the behavior of dilated convolutions. Hardswish activation is employed to harness heightened non-linear characteristics, enabling the network to extract more intricate acoustic features.The application of DSDCN is avoided in the input encoder and magnitude-phase decoder of the network. This strategic exclusion arises from the inherent flexibility of deformable convolutions in achieving adaptable receptive fields through feature offsets, as the use of DCN across the full spectrum of the input and output codec stages is not efficient.
@@ 2.4 Dense Convolution Codec
In the input-output encoder, we borrowed the Codec from MP-SEnet [11]. The Dilated Dense-Net structure comprises four dilated convolutional blocks with dense residual connections [23], each block having dilation factors set to ${1,2,4,8}$. In dilated convolutions, our primary objective is to augment the receptive field while preserving the kernel size and layer counts, ensuring the efficient capture of globally and intricately featured information.
@@ 3.2 Model setup
Throughout the training process, speech data was uniformly segmented into 30700 points, employing an FFT size of 510, a window length of 510, a hop length of 100, and a sampling rate of 16000111https://github.com/huaidanquede/MUSE-Speech-Enhancement. The training configuration encompassed a batch size of 2, a learning rate set to 0.0005 with a decay factor of 0.99, dense channels is 16, all the models were trained using the AdamW [30] optimizer until 100 epochs. During training, a single 8GB RTX 3070ti GPU was utilized.